OpenStudy (anonymous):

What is the simplified form of the expression 3[14 ÷ (42 – 32) – 5]?

OpenStudy (anonymous):

\[-\frac{ 54 }{ 5 }\]

OpenStudy (anonymous):

hang on im typing how to get it

OpenStudy (anonymous):

wait thats the wrong problem is suppose to be 3[14 ÷ (4 squared – 32 squared) – 5]?

OpenStudy (anonymous):

\[3(\left(\begin{matrix}14 \\ 42 - 32\end{matrix} -5\right)\] substract 32 from 42, gives you 10 \[3 \left(\begin{matrix}14 \\ 10\end{matrix} -5\right)\] in 14/10, the gcd is 2, so \[\frac{ 14 }{ 10 } = \frac{ 2x7 }{ 2x5 } = \frac{ 2 }{ 2 } x \frac{ 7 }{ 5 } = \frac{ 7 }{ 5 }\] \[3\left(\begin{matrix}7 \\ 5\end{matrix} -5\right)\] put \[\frac{ 7 }{ 5 } -5\] over the common denominator \[5. \frac{ 7 }{ 5 } + \frac{ 5(-5) }{ 5 } :\] \[3 \frac{ 7 }{ 5 } + \frac{ -5 x 5 }{ 5 } \] add the fractons over the common demnominator to a single fraction \[3 \frac{ 7 - 5x5 }{ 5 }\] mutiply -5 and 5 togather \[3 \frac{ -(25 -7) }{ 5 }\] subtract 7 from 25 \[3x \frac{ -18 }{ 5 }\] express it as a single fraction \[\frac{ -18x3 }{ 5 }\] multiply -18 and 3 togather gives you 54 \[-\frac{ 54 }{ 5 }\]

OpenStudy (anonymous):

oh god, lol ok hang on, just seen you typed it wrong

OpenStudy (anonymous):

\[-\frac{ 361 }{ 24 }\] for the updated problem, I think, im just starting the squares so I would have someone double check

OpenStudy (anonymous):

\[3 \left(\begin{matrix}14 \\ 4^2 -32^2\end {matrix}-5\right) \] if thats what it looks like then \[-\frac{ 361 }{ 24 }\] is the answer

OpenStudy (anonymous):

also as a mixed fraction it would be \[- 15\frac{ 1 }{ 24 }\]